Give the equation of the line perpendicular to 2x+3y=7 and passing through (6,4)
Answer by Alan3354(69443)

Give the equation of the line perpendicular to 2x+3y=7 and passing through (6,4)
-------------
Step 1, find the slope of the line 2x+3y=7
Step 2, find the slope of lines perpendicular.
(It's the negative inverse of the slope of 2x+3y=7)
Step 3, find the eqn with slope m thru (6,4)
y-4 = m*(x-6)